Solve: w^2 + 5w - 36 = 0
Answer by algebrahouse.com(1659)

w² + 5w - 36 = 0
(w + 9)(w - 4) = 0 {factored into two binomials}
w + 9 = 0 or w - 4 = 0 {set each factor equal to 0}
w = -9 or w = 4 {subtracted 9 and added 4, respectively}
For more help like this, ask the house at www.algebrahouse.com